Consider a risky portfolio. The end-of-year cash flow derived from the portfolio will be either $50,000 or $150,000, with equal probabilities of 0.5. The alternative riskless investment in T-bills pays 5%. a. If you require a risk premium of 10%, how much will you be willing to pay for the portfolio? (Round your answer to the nearest dollar amount.) b. Suppose the portfolio can be purchased for the amount you found in (a). What will the expected rate of return on the portfolio be? (Do not round intermediate calculations. Round your answer to the nearest whole percent.) c. Now suppose you require a risk premium of 15%. What is the price you will be willing to pay now? (Round your answer to the nearest dollar amount.)
Answer #1

a)

Your required rate of return on the portfolio= Risk Free Return+Risk Premium=5%+10%=15%

Expected value of the portfolio= 0.5*50000+0.5*150000=$100,000

Say, x is the amount you will be willing to pay for the portfolio.

Hence, x*(1+15%)=100000

or, x= $86956.52

You would be willing to pay $86956.52 for the portfolio.

b) Expected return on the portfolio= (100000-86956.52)/86956.52=15%

c) Similarly, if the risk premium is 15%, then required rate of return=5%+15%=20%

Hence, price you will be willing to pay= 100000/(1+20%)=$83333.33